UNTANGLING THE NEAR-IR HIGH RESOLUTION SPECTRUM OF THE METHYL PEROXY RADICAL

\underline{PATRICK DUPRÉ}, PATRICK RUPPER, SHENGHAI WU and TERRY A. MILLER, Department of Chemistry, Ohio State University, 120 W. 18th Ave, Columbus OH, 43210.

The jet-cooled high resolution CRDS spectra of the methyl peroxy radical, CH_{3}OO and CD_{3}OO has been obtained for the weak Ã^{2}A′ ← X̃^{2}A′′ 0_{0}^{0} band. The rotational spectra of the isotopologue species exhibit very different structure. The hindered internal rotational involving methyl and O_{2} groups is a source of this very different structure, due to the strong dependence of the tunneling splitting upon mass through the F parameter (∼ 4.4 cm^{−1} for CD_{3}OO compared to ∼ 7.0 cm^{−1} for CH_{3}OO). We will discuss the formalism in the limit of the high torsional barrier in the light of the experimental results, emphasizing the spectral resemblances and differences.